Assume the likelihood that any flight on Northwest Airlines arrives within 15 minutes of the scheduled time is 0.90. We select four flights from yesterday for study. (Question 50)
a. What is the likelihood all four of the selected flights arrived within 15 minutes of scheduled time?
b. What is the likelihood at least one of the selected flights did not arrive within 15 minutes of the scheduled time?
c. What is the likelihood that none of the selected flights arrived within 15 minutes of the scheduled time?

Answer by stanbon(75887)

Assume the likelihood that any flight on Northwest Airlines arrives within 15 minutes of the scheduled time is 0.90. We select four flights from yesterday for study. (Question 50)
a. What is the likelihood all four of the selected flights arrived within 15 minutes of scheduled time?
Ans: 0.90^4 = 0.6561
-----------------------------------
b. What is the likelihood at least one of the selected flights did not arrive within 15 minutes of the scheduled time?
Ans: 1 - (0.10^4) = 0.9999..
c. What is the likelihood that none of the selected flights arrived within 15 minutes of the scheduled time?
Ans: 0.10^4 = 0.0001
=========================
Cheers,
Stan H.